# 3D sans 3D Scans: Scalable Pre-training from Video-Generated Point Clouds

### v1

@@ Abstract
Despite recent progress in 3D self-supervised learning, collecting large-scale 3D scene scans remains expensive and labor-intensive. In this work, we investigate whether 3D representations can be learned from unlabeled videos recorded without any real 3D sensors. We present Laplacian-Aware Multi-level 3D Clustering with Sinkhorn-Knopp (LAM3C), a self-supervised framework that learns from video-generated point clouds from unlabeled videos. We first introduce RoomTours, a video-generated point cloud dataset constructed by collecting room-walkthrough videos from the web (e.g., real-estate tours) and generating 49,219 scenes using an off-the-shelf feed-forward reconstruction model. We also propose a noise-regularized loss that stabilizes representation learning by enforcing local geometric smoothness and ensuring feature stability under noisy point clouds. Remarkably, without using any real 3D scans, LAM3C achieves higher performance than the previous self-supervised methods on indoor semantic and instance segmentation. These results suggest that unlabeled videos represent an abundant source of data for 3D self-supervised learning.
@@ 1 Introduction
Understanding real-world 3D scenes is essential for AI systems that require visual spatial intelligence [yang2025thinking]. As of 2025, vision foundation models (VFMs) such as DINOv2/v3 [oquab2023dinov2, simeoni2025dinov3] and SAM [Kirillov_2023_ICCV] have shown remarkable generalization in image recognition tasks. DINOv3 is pre-trained on about 1.7 billion unlabeled images, while SAM is trained on 11 million densely annotated images. In contrast, 3D data is highly constrained by scanning real environments and manual annotation, making large-scale collection extremely difficult. The largest widely used indoor scene dataset offers only about 5k unique scenes [baruch2021arkitscenes]. This disparity in data scalability poses significant limitations to achieving scalable 3D pre-training.
@@
3D self-supervised learning (3D-SSL) for indoor scene understanding has made steady progress [wu2024towards, wu2023masked, xie2020pointcontrast]. For instance, Sonata [wu2025sonata] is a representative approach that extends DINOv2 to point clouds and mitigates the geometric shortcut problem, improving segmentation performance. However, most existing 3D-SSL methods remain fundamentally limited by the availability of 3D data. Even when combining real and synthetic point clouds, the training scale is around 140k samples, and merely 18k for real 3D scans alone. Such limited data scale hinders 3D-SSL methods from reaching the same level of success achieved in 2D vision. This limitation motivates us to develop a new 3D pre-training framework without relying on real 3D scans.
@@
We hypothesize that we can extract enough geometric cues from unlabeled videos to learn 3D representations. Traditionally, computer vision has aimed to infer 3D structures from multi-view images. Classical methods such as Structure-from-Motion (SfM) [ullman1979interpretation, wu2011visualsfm, agarwal2011building, tomasi1992shape] and Multi-View Stereo (MVS) [okutomi1993multiple, furukawa2009accurate, seitz1999photorealistic] reconstruct 3D structure by identifying correspondences across multi-view images and optimizing camera parameters and point clouds. While these classical methods are accurate, they involve a complex optimization procedure. More recently, feed-forward reconstruction models such as VGGT have emerged [wang2025vggt, Wang_2024_CVPR, keetha2025mapanything]. Despite their simple design, these models can directly infer 3D structure from multi-view images and achieve reconstructions that are comparable to, or even better than, classical methods. These advances indicate that videos contain strong geometric cues that modern reconstruction models can extract. Building on this insight, we argue that this opens a path towards scan-free indoor scene understanding.
@@
We aim to leverage video-generated point clouds reconstructed from unlabeled videos as pre-training data and show that they are remarkably effective pre-training signals for 3D-SSL. We introduce Laplacian-Aware Multi-level 3D Clustering with Sinkhorn-Knopp (LAM3C), a pre-training framework that learns stable representations from video-generated point clouds via Sinkhorn-Knopp clustering [knight2008sinkhorn, cuturi2013sinkhorn, asano2019self, caron2020unsupervised]. LAM3C comprises the following key components: (i) RoomTours. We construct a new dataset, RoomTours, by collecting room-walkthrough videos from the web (e.g., real-estate tours) and reconstructing video-generated point clouds from various camera locations using an off-the-shelf feed-forward 3D reconstruction model. RoomTours contains 49,219 scenes from video-generated point clouds across indoor videos. (ii) Noise-regularized loss. Video-generated point clouds contain noise and missing regions, which makes point-wise embeddings unstable during Sinkhorn-Knopp clustering. To address this, we introduce a noise-regularized loss composed of two terms: a Laplacian smoothing term, which encourages spatially adjacent points to produce similar embeddings by smoothing features along the local geometry of the video-generated point cloud; a noise consistency term, which enforces two views from the same scene to have similar global representations even in the presence of noisy regions. Together, these losses make learning robust to noisy video-generated point clouds and lead to stable representations. The main contributions of this paper are as follows.
@@
  • We propose a pipeline for generating large datasets of video-generated point clouds of indoor scenes from unlabeled videos. Thanks to this pipeline, we create RoomTours, a dataset of 49k point clouds obtained from videos collected from the web.
@@
  • We show that video-generated point clouds are sufficient to learn powerful representations. We actually outperform Sonata [wu2025sonata] trained only on real 3D scans on several benchmarks (see Fig. 1). On top of LAM3C, the keys are a sufficiently large dataset of video-generated point clouds, like RoomTours, a large-capacity 3D backbone, and sufficiently long pre-training schedule.
@@ 2 Related work
SSL on point clouds. Manual annotation of real 3D scans is costly, making 3D-SSL, which generates pseudo supervision from unlabeled data, an essential technique for representation learning. Early 3D-SSL methods concentrated on building representations for dense scans of single objects [chen2021shape, poursaeed2020self, sauder2019self]. In indoor scene understanding, PointContrast [xie2020pointcontrast] pioneered 3D-SSL by performing point-level contrastive learning, establishing the first successful pre-training on large-scale point clouds. Subsequent works improved performance thanks to a better choice of contrasting pairs [depthcontrast, strl, liu2023fac, wu2023masked, groupcontrast, Hou_2021_CVPR], and also extended the application to sparse point clouds of large outdoor scenes [proposalcontrast, segcontrast, tarl, bevcontrast]. Beyond contrastive methods, other strategies used a reconstruction pretext task [voxelmae1, voxelmae2, pointmae, pointm2ae, pointbert, also]. Recently, inspired by the success of DINO [caron2021emerging, oquab2023dinov2], Sonata adapted this technique to point clouds with recipes to mitigate low-level geometric shortcuts during pre-training. It achieved notable gains over the state-of-the-art, especially in linear probing.
@@
Feed-forward reconstruction models. Feed-forward reconstruction models have emerged, marking a significant shift in the paradigm of 3D reconstruction [Wang_2024_CVPR, Cabon_2025_CVPR, wang2025vggt, Wang_2024_CVPR, keetha2025mapanything]. VGGT [wang2025vggt] jointly predicts camera poses and point maps from multi-view images using a single transformer model, eliminating the need for sequential SfM optimization [okutomi1993multiple, furukawa2009accurate, seitz1999photorealistic]. In addition, $\pi^{3}$ [wang2025pi] introduces a fully permutation-equivariant architecture, enabling affine-invariant camera poses and scale-invariant local point maps without a reference view that is invariant to the ordering of input views. This design yields a more robust reconstruction that is well suited to in-the-wild domains. While these models achieve high-quality reconstruction without explicit geometric optimization, they only focus on reconstruction itself rather than representation learning for 3D scene understanding. The potential of the intermediate features or reconstructed point clouds for pre-training has not yet been fully realized.
This paper explores the potential of video-generated point clouds, produced by a feed-forward reconstruction model, as pre-trained data for understanding indoor scenes. Our proposed LAM3C achieves representation learning without real 3D scans.
@@ 3 LAM3C
Figure 2: Laplacian smoothing loss. LAM3C resolves learning instability by using kNN graphs for video-generated point clouds and assigning distance-based weights. This reduces embedding divergence caused by noisy point clouds and promotes learning more stable representations from neighboring point clouds.
This section introduces our proposed LAM3C as a pre-training framework. LAM3C learns 3D representations through Sinkhorn-Knopp clustering from video-generated point clouds generated by a feed-forward reconstruction model. The architecture follows a teacher–student design, where the student model is trained to mimic the feature representations produced by the teacher. The teacher model is updated using an exponential moving average (EMA) [polyak1992acceleration] of the student parameters. This design encourages more generalizable representations, whereas robustness to noisy or missing point clouds comes from our noise-regularized losses. In addition to the standard clustering loss [wu2025sonata], LAM3C introduces two new losses that mitigate the noisy nature of video-generated point clouds to enhance local smoothness and feature consistency. We recall the Sonata loss and describe our new noise-regularized losses below.
@@
Here, $p_{i}^{(s)}=\mathrm{softmax}(z_{i}^{(s)}/\tau_{s})$ denotes the student’s probability distribution, and $q_{i}^{(t)}=\mathrm{Sinkhorn}(z_{i}^{(t)}/\tau_{t})$ represents the teacher’s entropy-regularized soft assignment, following [asano2019self, caron2020unsupervised]. Both $z_{i}^{(s)}$ and $z_{i}^{(t)}$ are prototype logits, while $\tau_{s}$ and $\tau_{t}$ are temperature parameters that control the sharpness of the distributions.
Figure 3: Noise consistency loss. This is a constraint term stating that the same point cloud for the teacher and student models should yield the same embedding in the feature space. This enables more stable clustering even in video-generated point clouds.
Laplacian smoothing loss. Video-generated point clouds contain noise and missing regions due to imperfect reconstruction. To stabilize representation learning under such point clouds, we construct a k-Nearest Neighbor (kNN) graph on each point cloud and apply a Laplacian smoothness loss as shown in Fig. 2. For a point $x_{i}$ with embedding $z_{i}=f_{\theta}(x_{i})$, the loss encourages nearby points to produce similar embeddings:
@@
We assign each edge a distance-based weight, $w_{ij}=\exp(-{\|p_{i}-p_{j}\|^{2}}/{\sigma^{2}}),$ so that closer points contribute more to the smoothness constraint. We compute the edge weight $w_{ij}$ based on the $L2$ distance between points in the video-generated point cloud. The scale parameter $\sigma$ is adaptively estimated per point cloud as the median of the kNN distance, making the weighting robust to variations in point density. In addition, neighbors further away than a certain threshold are removed for improved robustness.
@@
Noise consistency loss. While Laplacian smoothing enforces geometric smoothness across neighboring points, our noise consistency loss ensures feature stability for different augmented views of the same point cloud. As shown in Fig. 3, given two views of the same video-generated point cloud $x^{(a)}$ and $x^{(b)}$, we feed them into the EMA teacher $g_{\text{EMA}}$ and the student $f_{\theta}$, respectively, and minimize the discrepancy between their embeddings:
@@
where $\mathcal{P}$ denotes the set of kNN correspondences between the noise-augmented global-view $x^{(a)}$ and the masked global-view $x^{(b)}$. Each pair $(i,j)\in\mathcal{P}$ represents two corresponding points from the teacher and student inputs, respectively. $g_{\text{EMA}}(\cdot)$ and $f_{\theta}(\cdot)$ denote the backbone encoders that output point-wise embeddings. Minimizing $R_{\text{cons}}$ enforces that the same point maintains a consistent embedding even in the presence of noise.
@@
During pre-training, we schedule the strength of regularization. The Laplacian coefficient $\lambda$ starts at $2\text{e}{-4}$ and linearly decays to $3\text{e}{-3}$, whereas the noise consistency coefficient $\mu$ is kept fixed at $0.05$. We describe the exploratory experiments regarding parameters $\lambda$ and $\mu$ in the supplementary material. By combining Laplacian smoothing and noise consistency, noise-regularized clustering enables stable representation learning, even from noisy point clouds.
@@ 4 RoomTours
This section details RoomTours, generating video-generated point clouds from unlabeled videos on the web.
@@
To ensure accurate 3D reconstruction, we prioritize videos with fewer dynamic objects (e.g., real-estate agents walking in front of the camera) and minimal editing cuts or aggressive camera motion. Once candidate channels are selected, videos are automatically filtered based on metadata such as duration and undesired keywords (e.g., “CG,” “drone,” “short”). Only videos that satisfy these criteria are downloaded, resulting in a total of 3,462 videos. In addition, we utilize videos from RealEstate10k [zhou2018stereo], the YouTube House Tours Dataset [chang2020semantic], and HouseTours [Celen_2025_ICCV].
Figure 4: Overview of RoomTours construction. We segment the video into scene sequences using CLIP [radford2021learning], and generate video-generated point clouds by inputting each scene sequence into $\pi^{3}$. Because the scenes differ from real 3D scans in coordinate system, scale, and spacing, we apply a post-processing alignment.
Figure 5: Visualization of video-generated point clouds from the RoomTours. These are pseudo-scenes generated as video-generated point clouds from unsupervised videos collected from the web. Visually, they achieve very high-quality reconstruction of real-world indoor scenes. However, for example, the leftmost scene contains a large amount of noisy point clouds in the scene due to camera shake in the input video, resulting in blurred object boundaries and cases where walls or floors appear doubled.
@@
We apply frame-level zero-shot scene classification using CLIP [radford2021learning]. In the first step, each frame is classified as indoor or outdoor, and frames predicted as outdoor are discarded. In the second step, the remaining indoor frames are grouped into three broad scene types, namely “living room”, “bedroom”, and “bathroom”. These categories are deliberately coarse because fine categories would split a video into many short sequences, making reconstruction unreliable and more coarse categories would mix widely different scenes into a single sequence. Scene boundaries are detected by changes in the predicted scene type, and the video is split accordingly. We stabilize the predictions by enforcing temporal consistency, ensuring the same class label within 0.5 seconds.
Video-generated point clouds generation. We create video-generated point clouds from videos using $\pi^{3}$ [wang2025pi], which is a feed‑forward 3D reconstruction model that does not rely on SLAM [cheeseman1987stochastic, davison2007monoslam] and Bundle Adjustment [triggs1999bundle] optimization. Given a scene sequence, we first sample frames at uniform intervals and, for long sequences, automatically downsample to a target number of frames (e.g., 200-400 frames per scene). Each frame is resized to meet a per-image pixel count while preserving aspect ratio. $\pi^{3}$ processes the sequence in a single mixed-precision forward pass. It filters unreliable predictions using confidence masking and edge suppression, removes remaining outliers, and then fuses the per-frame estimates into final colored point clouds. Our method outputs colored point clouds for each scene, with an average runtime of about 5 minutes. Example outputs of RoomTours are shown in Fig. 5.
Table 1: Indoor Semantic Segmentation Results across Datasets. All self-supervised methods are evaluated by full fine-tuning (Full-FT) or linear probing (LP) on PTv3 (Base) for 100 epochs. We use mIoU as evaluation metric. VGPC indicates video-generated point clouds.
Semantic Seg. Pretraining Data ScanNet [dai2017scannet] ScanNet200 [rozenberszki2022language] ScanNet++ Val [yeshwanth2023scannet++] S3DIS Area 5 [Armeni_2016_CVPR]
@@
[…]
* uses PTv3 (Large) and 434k pre-training steps. gray for sup. method or access to val/test splits at pre-training.
Table 2: Indoor Instance Segmentation Results across Datasets. All self-supervised methods are evaluated by full fine-tuning (Full-FT) or linear probing (LP) on PTv3 (Base) for 100 epochs. We use mAP as evaluation metric. VGPC indicates video-generated point clouds.
Instance Seg. Pretraining Data ScanNet [dai2017scannet] ScanNet200 [rozenberszki2022language] ScanNet++ Val [yeshwanth2023scannet++] S3DIS Area 5 [Armeni_2016_CVPR]
@@
[…]
* uses PTv3 (Large) and 434k pre-training steps. gray for sup. method or access to val/test splits at pre-training.
Alignment of scenes from video-generated point clouds. A scene reconstructed from $\pi^{3}$ often contains noise and has inconsistent coordinate system or scale. Thus, we apply a lightweight alignment procedure that enforces geometric consistency while maintaining data diversity.
First, we randomly sample points from video-generated point clouds and remove statistical outliers using a standard SOR filter, which eliminates points with abnormally large k-NN distances. Next, we detect a dominant plane and align the scene to a Z-up coordinate system. We then perform an SVD refinement of its orientation. If plane detection fails, ceilings or walls may be mistakenly selected, causing the coordinate system to flip. We keep such scenes unchanged in our dataset because automatic rejection could remove valid samples and reduce diversity.
In addition, we define the scale of a scene as the diagonal length of its axis-aligned bounding box, which captures the overall spatial extent of the geometry. This value is denoted as $s_{\text{current}}$. We then align the scene with a factor $\alpha=\frac{s_{\text{target}}}{s_{\text{current}}}$, where $s_{\text{target}}$ is drawn from a scale distribution estimated in advance over ScanNet.
Finally, we estimate per-point normals using local PCA. This process aligns video-generated point clouds, converting them into a geometrically consistent pre-training data.
Dataset statistics. The final RoomTours contains 49,219 scenes from video-generated point clouds, covering diverse lighting conditions, camera trajectories, and architectural styles across different countries. Compared to existing 3D scans, RoomTours captures more diversity 3D structure, providing a scalable foundation for learning stable representations from unlabeled videos. Here, we denote the full version, which contains 49,219 video-generated point clouds including videos from existing datasets, as RoomTours-49k (or LAM3C-49k for the model trained on RoomTours-49k). We also refer to the subset consisting of 15,921 video-generated point clouds built solely from our collected videos as RoomTours-16k (or LAM3C-16k for the model trained on RoomTours-16k). We provide the dataset details in the supplementary materials.

### v2

@@ Abstract
Despite recent progress in 3D self-supervised learning, collecting large-scale 3D scene scans remains expensive and labor-intensive. In this work, we investigate whether 3D representations can be learned from unlabeled videos recorded without any real 3D sensors. We present Laplacian-Aware Multi-level 3D Clustering with Sinkhorn-Knopp (LAM3C), a self-supervised framework that learns from video-generated point clouds reconstructed from unlabeled videos. We first introduce RoomTours, a video-generated point cloud dataset constructed by collecting room-walkthrough videos from the web (e.g., real-estate tours) and generating 49,219 scenes using an off-the-shelf feed-forward reconstruction model. We also propose a noise-regularized loss that stabilizes representation learning by enforcing local geometric smoothness and ensuring feature stability under noisy point clouds. Remarkably, without using any real 3D scans, LAM3C achieves better performance than previous self-supervised methods on indoor semantic and instance segmentation. These results suggest that unlabeled videos represent an abundant source of data for 3D self-supervised learning. Our source code is available at https://github.com/ryosuke-yamada/lam3c.
@@ 1 Introduction
Understanding real-world 3D scenes is essential for AI systems that require visual spatial intelligence [55]. As of 2025, vision foundation models (VFMs) such as DINOv2/v3 [30, 39] and SAM [22] have shown remarkable generalization in image recognition tasks. DINOv3 is pre-trained on about 1.7 billion unlabeled images, while SAM is trained on 11 million densely annotated images. In contrast, 3D data is highly constrained by scanning real environments and manual annotation, making large-scale collection extremely difficult. The largest widely used indoor scene dataset offers only about 5k unique scenes [4]. This disparity in data scalability poses significant limitations to achieving scalable 3D pre-training.
@@
3D self-supervised learning (3D-SSL) for indoor scene understanding has made steady progress [52, 53, 54]. For instance, Sonata [49] is a representative approach that extends DINOv2 to point clouds and mitigates the geometric shortcut problem, improving segmentation performance. However, most existing 3D-SSL methods remain fundamentally limited by the availability of 3D data. Even when combining real and synthetic point clouds, the training scale is around 140k samples, and merely 18k for real 3D scans alone. Such limited data scale hinders 3D-SSL methods from reaching the same level of success achieved in 2D vision. This limitation motivates us to develop a new 3D pre-training framework without relying on real 3D scans.
@@
We hypothesize that we can extract enough geometric cues from unlabeled videos to learn 3D representations. Traditionally, computer vision has aimed to infer 3D structures from multi-view images. Classical methods such as Structure-from-Motion (SfM) [43, 48, 1, 41] and Multi-View Stereo (MVS) [29, 17, 38] reconstruct 3D structure by identifying correspondences across multi-view images and optimizing camera parameters and point clouds. While these classical methods are accurate, they involve a complex optimization procedure. More recently, feed-forward reconstruction models such as VGGT have emerged [45, 46, 21]. Despite their simple design, these models can directly infer 3D structure from multi-view images and achieve reconstructions that are comparable to, or even better than, classical methods. These advances indicate that videos contain strong geometric cues that modern reconstruction models can extract. Building on this insight, we argue that this opens a path towards scan-free indoor scene understanding.
@@
We aim to leverage video-generated point clouds (VGPC) reconstructed from unlabeled videos as pre-training data and show that they are remarkably effective pre-training signals for 3D-SSL. We introduce Laplacian-Aware Multi-level 3D Clustering with Sinkhorn-Knopp (LAM3C), a pre-training framework that learns stable representations from VGPC via Sinkhorn-Knopp clustering [23, 14, 3, 7]. LAM3C comprises the following key components: (i) RoomTours. We construct a new dataset, RoomTours, by collecting room-walkthrough videos from the web (e.g., real-estate tours) and reconstructing VGPC from various camera locations using an off-the-shelf feed-forward 3D reconstruction model. RoomTours contains 49,219 VGPC scenes reconstructed from indoor videos. (ii) Noise-regularized loss. VGPC contain noise and missing regions, which makes point-wise embeddings unstable during Sinkhorn-Knopp clustering. To address this, we introduce a noise-regularized loss composed of two terms: a Laplacian smoothing term, which encourages spatially adjacent points to produce similar embeddings by smoothing features along the local geometry of the VGPC; a noise consistency term, which enforces two views from the same scene to have similar global representations even in the presence of noisy regions. Together, these losses make learning robust to noisy VGPC and lead to stable representations. The main contributions of this paper are as follows.
@@
  • We propose a pipeline for generating large datasets of VGPC of indoor scenes from unlabeled videos. Thanks to this pipeline, we create RoomTours, a dataset of 49k point clouds obtained from videos collected from the web.
@@
  • We show that VGPC are sufficient to learn 3D representations. We outperform Sonata [49] trained only on real 3D scans on several benchmarks (see Fig. 1). Beyond LAM3C, the key ingredients are a sufficiently large dataset of VGPC, like RoomTours, a large-capacity 3D backbone, and sufficiently long pre-training schedule.
@@ 2 Related work
SSL on point clouds. Manual annotation of real 3D scans is costly, making 3D-SSL, which generates pseudo-supervision from unlabeled data, an essential technique for representation learning. Early 3D-SSL methods concentrated on building representations for dense scans of single objects [12, 33, 36]. In indoor scene understanding, PointContrast [54] pioneered 3D-SSL by performing point-level contrastive learning, establishing the first successful pre-training on large-scale point clouds. Subsequent works improved performance thanks to a better choice of contrasting pairs [61, 20, 24, 53, 44, 19], and also extended the application to sparse point clouds of large outdoor scenes [57, 27, 28, 37]. Beyond contrastive methods, other strategies used a reconstruction pretext task [26, 18, 31, 59, 58, 5]. Recently, inspired by the success of DINO [8, 30], Sonata adapted this technique to point clouds with recipes to mitigate low-level geometric shortcuts during pre-training. It achieved notable gains over the state-of-the-art, especially in linear probing.
@@
Feed-forward reconstruction models. Feed-forward reconstruction models have emerged, marking a significant shift in the paradigm of 3D reconstruction [46, 6, 45, 21]. VGGT [45] jointly predicts camera poses and point maps from multi-view images using a single transformer model, eliminating the need for sequential SfM optimization [29, 17, 38]. In addition, $\pi^{3}$ [47] introduces a fully permutation-equivariant architecture, enabling affine-invariant camera poses and scale-invariant local point maps without a reference view that is invariant to the ordering of input views. This design yields a more robust reconstruction that is well suited to in-the-wild domains. While these models achieve high-quality reconstruction without explicit geometric optimization, they only focus on reconstruction itself rather than representation learning for 3D scene understanding. The potential of the intermediate features or reconstructed point clouds for pre-training has not yet been fully realized.
This paper explores the potential of VGPC, produced by a feed-forward reconstruction model, as pre-training data for understanding indoor scenes. Our proposed LAM3C achieves representation learning without real 3D scans.
@@ 3 LAM3C
Figure 2: Laplacian smoothing loss. LAM3C resolves learning instability by using kNN graphs for VGPC and assigning distance-based weights. This reduces embedding divergence caused by noisy point clouds and promotes more stable representations from neighboring point clouds.
This section introduces our proposed LAM3C as a pre-training framework. LAM3C learns 3D representations through Sinkhorn-Knopp clustering from VGPC generated by a feed-forward reconstruction model. The architecture follows a teacher–student design, where the student model is trained to mimic the feature representations produced by the teacher. The teacher model is updated using an exponential moving average (EMA) [32] of the student parameters. This design encourages more generalizable representations, whereas robustness to noisy or missing point clouds comes from our noise-regularized losses. In addition to the standard clustering loss [49], LAM3C introduces two new losses that mitigate the noisy nature of VGPC to enhance local smoothness and feature consistency. We recall the Sonata loss and describe our new noise-regularized losses below.
@@
Here, $p_{i}^{(s)}=\mathrm{softmax}(z_{i}^{(s)}/\tau_{s})$ denotes the student’s probability distribution, and $q_{i}^{(t)}=\mathrm{Sinkhorn}(z_{i}^{(t)}/\tau_{t})$ represents the teacher’s entropy-regularized soft assignment, following [3, 7]. Both $z_{i}^{(s)}$ and $z_{i}^{(t)}$ are prototype logits, while $\tau_{s}$ and $\tau_{t}$ are temperature parameters that control the sharpness of the distributions.
Figure 3: Noise consistency loss. This is a constraint term stating that the same point cloud for the teacher and student models should yield the same embedding in the feature space. This enables more stable clustering even in VGPC.
Laplacian smoothing loss. VGPC contain noise and missing regions due to imperfect reconstruction. To stabilize representation learning under such point clouds, we construct a k-Nearest Neighbor (kNN) graph on each point cloud and apply a Laplacian smoothness loss as shown in Fig. 2. For a point $x_{i}$ with embedding $z_{i}=f_{\theta}(x_{i})$, the loss encourages nearby points to produce similar embeddings:
@@
We assign each edge a distance-based weight, $w_{ij}=\exp(-{\|p_{i}-p_{j}\|^{2}}/{\sigma^{2}}),$ so that closer points contribute more to the smoothness constraint. We compute the edge weight $w_{ij}$ based on the $L2$ distance between points in the VGPC. The scale parameter $\sigma$ is adaptively estimated per point cloud as the median of the kNN distance, making the weighting robust to variations in point density. In addition, neighbors further away than a certain threshold are removed for improved robustness. In practice, we compute the regularization on a kNN graph with distance-based edge weights and degree normalization. To improve robustness to noisy reconstructions and outliers, we replace the squared $L2$ edge penalty with a Huber penalty.
@@
Noise consistency loss. While Laplacian smoothing enforces geometric smoothness across neighboring points, our noise consistency loss ensures feature stability for different augmented views of the same point cloud. As shown in Fig. 3, given two views of the same VGPC $x^{(a)}$ and $x^{(b)}$, we feed them into the EMA teacher $g_{\text{EMA}}$ and the student $f_{\theta}$, respectively, and minimize the discrepancy between their embeddings:
@@
where $\mathcal{P}$ denotes the set of kNN correspondences between the noise-augmented global-view $x^{(a)}$ and the masked global-view $x^{(b)}$. Each pair $(i,j)\in\mathcal{P}$ represents corresponding points from the student and teacher inputs, respectively. $g_{\text{EMA}}(\cdot)$ and $f_{\theta}(\cdot)$ denote the backbone encoders that output point-wise embeddings. Minimizing $R_{\text{cons}}$ enforces that the same point maintains a consistent embedding even in the presence of noise.
@@
During pre-training, we schedule the strength of regularization. The Laplacian coefficient $\lambda$ starts at $2\text{e}{-4}$ and linearly increases to $3\text{e}{-3}$, while the noise consistency coefficient $\mu$ is fixed at $0.05$. Additional ablations on $\lambda$ and $\mu$ are provided in the supplementary material. By combining Laplacian smoothing with noise consistency, our noise-regularized clustering enables stable representation learning even from noisy point clouds.
LAM3C’s loss does not rely on hand-crafted indoor priors such as object scale or topology. In $L_{\text{clustering}}$, the student predicts targets computed from global views while observing only local views or masked global views. Since the targets encode global context, the task cannot be solved using local geometry alone, even for similar local structures. This encourages the representation to capture global context without relying on indoor priors.
@@ 4 RoomTours
This section describes RoomTours and the process used to generate VGPC from unlabeled videos on the web.
@@
To ensure accurate 3D reconstruction, we prioritize videos with fewer dynamic objects (e.g., real-estate agents walking in front of the camera) and minimal editing cuts or aggressive camera motion. Once candidate channels are selected, videos are automatically filtered based on metadata such as duration and undesired keywords (e.g., “CG,” “drone,” “short”). Only videos that satisfy these criteria are downloaded, resulting in a total of 3,462 videos. In addition, we utilize videos from RealEstate10k [62], the YouTube House Tours Dataset [10], and HouseTours [9].
Figure 4: Overview of RoomTours construction. We segment the video into scene sequences using CLIP [34], and generate VGPC by inputting each scene sequence into $\pi^{3}$. Because the scenes differ from real 3D scans in coordinate system, scale, and spacing, we apply a post-processing alignment.
Figure 5: Visualization of VGPC from the RoomTours. These are pseudo-scenes generated as VGPC from unsupervised videos collected from the web. Visually, they achieve very high-quality reconstruction of real-world indoor scenes. However, for example, the leftmost scene contains a large amount of noisy point clouds in the scene due to camera shake in the input video, resulting in blurred object boundaries and cases where walls or floors appear doubled.
@@
We apply frame-level zero-shot scene classification using CLIP [34]. In the first step, each frame is classified as indoor or outdoor, and frames predicted as outdoor are discarded. In the second step, the remaining indoor frames are grouped into three broad scene types, namely “living room”, “bedroom”, and “bathroom”. These categories are deliberately coarse because fine-grained categories would split a video into many short sequences, making reconstruction unreliable, whereas coarser categories would mix widely different scenes into a single sequence. Scene boundaries are detected by changes in the predicted scene type, and the video is split accordingly. We stabilize the predictions by enforcing temporal consistency, ensuring the same class label for frames within 0.5 seconds.
VGPC generation. We create VGPC from videos using $\pi^{3}$ [47], which is a feed‑forward 3D reconstruction model that does not rely on SLAM [11, 16] and Bundle Adjustment [42] optimization. Given a scene sequence, we first sample frames at uniform intervals and, for long sequences, automatically downsample to a target number of frames (e.g., 200-400 frames per scene). Each frame is resized to meet a per-image pixel count while preserving aspect ratio. $\pi^{3}$ processes the sequence in a single mixed-precision forward pass. It filters unreliable predictions using confidence masking and edge suppression, removes remaining outliers, and then fuses the per-frame estimates into final colored point clouds. Our method outputs colored point clouds for each scene, with an average runtime of about 5 minutes. Example outputs of RoomTours are shown in Fig. 5.
Table 1: Indoor Semantic Segmentation Results across Datasets. All methods are evaluated by full fine-tuning (Full-FT) or linear probing (LP) on PTv3 (Base) for 100 epochs. We use mIoU as evaluation metric. VGPC indicates video-generated point clouds.
Semantic Seg. Pre-training Data ScanNet [15] ScanNet200 [35] ScanNet++ Val [56] S3DIS Area 5 [2]
@@
* uses PTv3 (Large) and 437k pre-training steps. Gray indicates sup. method or access to val/test splits during pre-training.
Table 2: Indoor Instance Segmentation Results across Datasets. All methods are evaluated by full fine-tuning (Full-FT) or linear probing (LP) on PTv3 (Base) for 100 epochs. We use mAP as evaluation metric. VGPC indicates video-generated point clouds.
Instance Seg. Pre-training Data ScanNet [15] ScanNet200 [35] ScanNet++ Val [56] S3DIS Area 5 [2]
@@
* uses PTv3 (Large) and 437k pre-training steps. Gray indicates sup. method or access to val/test splits during pre-training.
Alignment of scenes from VGPC. A scene reconstructed from $\pi^{3}$ often contains noise and has an inconsistent coordinate system or scale. Thus, we apply a lightweight alignment procedure that enforces geometric consistency while maintaining data diversity. First, we randomly sample points from VGPC and remove statistical outliers using a standard SOR filter, which eliminates points with abnormally large k-NN distances. Next, we detect a dominant plane and align the scene to a Z-up coordinate system. We then perform an SVD refinement of its orientation. If plane detection fails, ceilings or walls may be mistakenly selected, causing the coordinate system to flip. We keep such scenes unchanged in our dataset because automatic rejection could remove valid samples and reduce diversity.
In addition, we define the scale of a scene as the diagonal length of its axis-aligned bounding box, which captures the overall spatial extent of the geometry. This value is denoted as $s_{\text{current}}$. We then align the scene with a factor $\alpha=\frac{s_{\text{target}}}{s_{\text{current}}}$, where $s_{\text{target}}$ is drawn from a scale distribution estimated in advance over ScanNet. Finally, we estimate per-point normals using local PCA. This process aligns VGPC, converting them into geometrically consistent pre-training data.
Dataset statistics. The final RoomTours contains 49,219 scenes from VGPC, covering diverse lighting conditions, camera trajectories, and architectural styles across different countries. Compared to existing 3D datasets, RoomTours captures a greater diversity of 3D structures, providing a scalable foundation for learning stable representations from unlabeled videos. Here, we denote the full version, which contains 49,219 VGPC including videos from existing datasets, as RoomTours-49k (or LAM3C-49k for the model trained on RoomTours-49k). We also refer to the subset consisting of 15,921 VGPC built solely from our collected videos as RoomTours-16k (or LAM3C-16k for the model trained on RoomTours-16k). We provide the dataset details in the supplementary materials.